# Modeling Language Evolution Using a Spin Glass Approach

### v1

@@ IV Introducing a Hebbian term
The model considered so far ignores the evidence provided by existing natural languages, in setting the random couplings - they are assigned purely at random. It may be hypothesized, however, that if specific realizations of syntax have arisen in human history, this may reflect interactions about particular syntactic parameters as they are implemented in the human brain. One simple-minded approach to delve into this complex and ill-defined issue is to introduce a set of Hebbian autoassociative synaptic weights, which in the syntax model would simply be additional pairwise interaction terms, strictly symmetric, that tend to stabilize existing realizations of language parameters. Inspired by the Hopfield model of an autoassociative network [27], we may identify the 58 languages in the PCM database as the ‘memories’ of the syntactic network. The new weights, denoted as $W_{ij}^{\text{Hop}}$, to be included in the interaction term in Eq. 1, $J_{ij\dots k}=W_{ij}^{\text{RND}}+\zeta I_{ij\dots k}+\rho W_{ij}^{\text{Hop}},$ are written as
@@
Figure 4 presents a schematic phase diagram in the $\zeta$-$\rho$ plane for a maximally fluid system – one with fully antisymmetric random interactions ($\phi=90^{\circ}$) – in the asymptotic regime (at the $100^{th}$ time step). Whatever the strength of the implications, there is clearly a cross-over from a fluid to a stable regime once the Hebbian term becomes effective, $\rho\gtrsim 1$: the Hebbian term then stabilizes the system even when the other interactions are strictly antisymmetric.
Figure 5 shows a similar phase diagram in the $\phi$-$\rho$ plane, when implications are negligible, $\zeta=0.01$ (note that $\rho$ has been moved to the y-axis). It appears to demonstrate that the stable regime, when syntactic parameters have completely stopped evolving by the $100^{th}$ time step, can be realized with either of two conditions, or both: when the Hebbian term is strong, $\rho>1$, or random interactions are sufficiently symmetric, $\phi<30^{\circ}$. The regime boundaries appear surprisingly close to a Boolean rule.
@@
Figure 6: Different languages rapidly converge, for $\rho>1$. Mean pairwise Hamming distances between realizations of distinct languages subject to identical random couplings, after only two time steps (($t=2$, panel (a)), and in the asymptotic regime ($t=100$, panel (b)). These distances vanish when the Hebbian term is dominant ($\rho>1$), indicating rapid collapse, already after 2 time steps, onto a single syntactic structure. Note that in the fluid regime, bottom right corners, Hamming distances $\sim 47$ imply that typically half the parameters agree, between any two vectors, and half do not.
In the remainder of this study, therefore, the Hebbian term will not be considered further, aside from a detail in Fig. 9 and for the analysis of syntactic distances presented in the Supplementary Material.
@@
Figure 8: Evolution of inter-language and intra-language distances. The diagrams, in the $\zeta$$\phi$ plane (at $\rho=0$), show both at early ($t=2$, top row) and asymptotic times ($t=100$, bottom row) the average Hamming distance between simulations starting from the same natural language (distance in condition S, left) and different natural languages (distance in condition D, middle), as well as their difference, (D-S, right), when using always identical random couplings. With time, both average distances diminish in areas characterized by strong implicational dynamics, especially when asymmetry is minimal. Still, they do not vanish: the average Hamming distance D between descendants of different languages tends (asymptotically, or for $t=100$) to decrease from D $\sim 35-37$ for weak implications ($\zeta<1$) thorough an intermediate range (e.g., D $\sim 18$ for $\zeta=10$, the star symbol) down to values D $\sim 3.1$ when implications dominate, for $\zeta>100$ (round symbol). Their average difference D-S illustrates the gradual fading of the phylogenetic/historical signal, i.e., the trace of shared ancestry among languages, everywhere except in the nearly symmetric and low (or at most intermediate) implicational strength regime, bottom left. Even there, the signal is however weak, a difference in at most $4-5$ parameters, compared to the dispersion observed starting from the same initial condition (left column).
@@ V Preservation of a phylogenetic signal
As an example of temporal evolution of a syntactic system within our model, we compare the presumed asymptotic state at $t=100$ with selected intermediate stages ($t=2,10,40$) for the case of a real-world starting point — Italian — in Fig. 7. When implications are irrelevant and asymmetry is high, the vector at $t=100$ remains very distant from its ancestors at essentially all preceding $t$ stages – note that a distance $\sim 47$ implies that half the 94 parameters take the same value in the two configurations, half a different value: effectively no syntax correlation. In other words, parameter values wander randomly in syntactic space, and reach no stable state nor do they approach an asymptotic state. Something similar occurs, albeit with reduced fluidity, when implications are stronger, but random couplings are still asymmetric. In contrast, the distance from the asymptotic configuration gradually vanishes if random couplings are prevailingly symmetric, and in fact when they are close to full symmetry syntactic parameters are already stable after a few time steps. These results are consistent across different languages.
@@
If that is the time course of simulations starting from the same point, we need to compare them with those starting from other initial syntax vectors. A phylogenetic signal encodes the relatedness among descendants of the same syntactic configuration that persists, despite the initially rambling trajectories in syntactic space, produced by the random couplings. We obviously need to keep the assignment of random couplings the same, and compare distances between different runs with the Same starting point (case S), and with Different starting points (case D). Figure 8 presents this comparison, as a grand average, in the entire $\zeta$-$\phi$ plane (at $\rho=0$), for configurations evaluated at two distinct time points: $t=2$ (top row) and $t=100$ (bottom row), corresponding to the early and asymptotic phases of the model, respectively. In each case, the main factor affecting these distances, which we label simply S and D like the cases they refer to, is the strength of the implications: when they are strong, they constrain the dynamics and both types of distance grow less - in fact, they decrease, reaching lower values at $t=100$ than at $t=2$. Thus, while for minimal asymmetry and in the low implication region, $\zeta<1$, the average D $\gtrsim 35$, in the strong implication regime, $\zeta>100$, it goes down below one tenth of this value, D $\sim 3.1$. When implication compete with the random couplings, $1001<\zeta<$ 100, one observes intermediate values; but, whatever $\zeta$, there is always dispersal, i.e., a multiplicity of trajectories, which do not converge to a single language.
@@
Interestingly, however, when we take the difference S-D, the only region where it remains above zero at $t=100$ (even though with small values) is the region of nearly symmetric random couplings and weak or intermediate implications, up to about $\zeta\approx 50$. Strong implications, $\zeta>100$, even with symmetric random couplings, retain a phylogenetic signal only for the first few time steps (top left region in the top right panel of Fig. 8). This indicates that in this model the phylogenetic signal is washed away, after a sufficient number of time steps, if implications are strong.
@@
Figure 9: Temporal dynamics of the phylogenetic signal. (a) The time course of the phylogenetic signal with $\phi$ fixed at $15^{\circ}$ and $\rho=0$, showing the rapid decrease in signal followed by stabilization at a steady-state value. A robust asymptotic signal is still observed for $\zeta=1$, i.e., when implications and random contributions are balanced. As $\zeta$ increases, indicating a greater influence of implications over random terms, the steady-state signal decreases. (b) The impact of varying $\rho$ and $\phi$, while keeping $\zeta=10$ constant. With limited asymmetry, $\phi=15^{\circ}$, a small Hebbian term (the red dashed line for $\rho=0.1$) leads the system to approximate the steady-state signal faster (compare to the yellow line for $\rho=0$). Further asymmetry, i.e., increasing $\phi$ leads to the vanishing of the signal, concurrent with the cross over to the fluid regime.
@@
VI Conclusions
@@
This study presents a framework for modeling syntactic evolution through the lens of disordered systems, bridging formal linguistics and statistical physics. By treating syntactic parameters as binary variables subject to both logical and typological implications and to random interactions, the model aims to capture the dual behavior of syntactic systems: their remarkable persistence across time and their capacity for gradual yet unpredictable diversification.
@@
Our results indeed indicate that in a portion of its phase space the network exhibits glassy dynamics, slow and gradually diverging, similar to that of spin glasses. Whether in the $N\to\infty$ limit the glassy regime would become a bona fide spin glass phase remains to be seen, but the glassy behavior is evident already in the network of limited size that models the $N=94$ syntactic parameters of the PCM database. Thus, the model suggests that syntactic evolution may not merely be the product of deterministic or functionally driven processes, and might instead be shaped by subtle, usually unobservable interactions that likely arise from the way syntax is realized in the brain of the speakers. The critical asymmetry value $\phi_{c}=30^{\circ}$, presumably marking a phase transition in the similar statistical physics model, can serve as a valid reference point for the syntax model as well, as it is not much affected by other factors, in particular the strength $\zeta$ of the implications. The suggestion from the model, then, is that natural language parameters interact, perhaps due to their implementation in the brain, via couplings that are, if not strictly symmetric, at least prevailingly symmetric (quantitatively, $\phi<30^{\circ}$).
@@
Of course, in our idealized framework – no external inputs whatsoever and zero ‘temperature’, i.e., zero noise – syntactic systems can become ‘frozen’ in certain configurations, while in more realistic scenarios they would continue to evolve, though slowly, in the corresponding region of the phase diagram. It should be noted that a separate study has shown that a disordered network of binary variables interacting randomly with each other as well as with a set of multi-state variables (also randomly) can be even glassier, that is slower, than if it were isolated [28]. The suggestion would then be that if syntax – the binary parameters – interacts randomly with e.g. the lexicon, or phonology, or semantics – which can all be formulated in terms of multi-state variables – it might evolve even slower than on its own.
@@
The introduction of a Hebbian memory term would stabilize syntactic configurations, even with asymmetric random couplings. However, an effective Hebbian term ($\rho>1$) would lead all languages to collapse onto the same syntax, at least in our implementation of the Hebbian term. More sophisticated versions of the same idea might avoid the collapse. For example, it might be argued that treating closely related languages, like several of those from the Indo-European family that comprises nearly half of the current PCM database, as adding their separate Hebbian contributions might be excessive, theoretically dubious, and might be the key factor leading to the collapse. Still, it is not obvious at this stage how to construct a ‘balanced’ Hebbian term, mimicking the effect of orthogonal memories in autoassociative memory networks [27].
Understanding better the conditions favoring glassy dynamics is also made urgent by the observation that in the present model, even in the absence of a Hebbian term, the phylogenetic signal rapidly vanishes if implications are sufficiently strong to really act as constraints, $\zeta\gtrsim 10-50$ (see Fig. 8); therefore invalidating the original assumption that adding weak interactions $W_{ij}$ to strong implications would suffice to lead to both divergence (the multiplicity of metastable states) and persistence (the phylogenetic signal). In the model as presented here, even with no Hebbian term ($\rho=0$), and with prevailingly symmetric random couplings ($\phi<30^{\circ}$, better if $\phi\sim 0^{\circ}$) divergence is maintained, to variable degrees, for all values of $\zeta$, but persistence requires weak or at most intermediate implications $\zeta\lesssim 50$.
This study, in conclusion, aims to present a first version of the mapping of syntactic parameters onto a recurrent network model, and needs to be extended in several directions before it can lead to well-grounded inferences. One direction we see as critical is a refined modeling of the undefined state of parameters, the definition of which has been disabled by the values taken by other parameters. Further, we should consider distinguishing different parameters, depending on the nature of their putative default state [29]. We plan to pursue these directions in future work.
Acknowledgments
@@
This study was started with an important initial contribution by Kwang Il Ryom, then a PhD student at SISSA, who should have coauthored it, had it been possible to contact him in recent months. We are also indebted to the many researchers who have contributed to the PCM database, foremost Cristina Guardiano and Paola Crisma; and to Eric De Giuli for stimulating discussions.
Data Availability
@@
The PCM Database has been published, e.g. in [9] and expanded versions are likely to be published in the near future. The software used in the present study will be made freely available upon publication.
Author Contributions
@@
H.Y. carried out the numerical experiments and drafted the manuscript; G.L. conceptualized the study, structured the syntactic implications in the PCM database and revised the manuscript; A.T. conceptualized the study, coded the implications as polysynaptic weights and drafted the manuscript.
Appendix A Dynamical roles of syntactic dependence classes

### v2

@@ IV Introducing a Hebbian term
The model considered so far ignores the evidence provided by existing natural languages, in setting the hidden couplings - they are assigned purely at random. It may be hypothesized, however, that if specific realizations of syntax have arisen in human history, this may reflect interactions about particular syntactic parameters as they are implemented in the human brain. One simple-minded approach to delve into this complex and ill-defined issue is to introduce a set of Hebbian autoassociative synaptic weights, which in the syntax model would simply be additional pairwise interaction terms, strictly symmetric, that tend to stabilize existing realizations of language parameters. Inspired by the Hopfield model of an autoassociative network [hopfield1982neural], which is known to preserve its glassy characteristics with moderate asymmetry [treves1988metastable], we may identify the 58 languages in the PCM database as the ‘memories’ of the syntactic network. The new weights, denoted as $W_{ij}^{\text{Hop}}$, to be included in the interaction term in Eq. 1, $J_{ij\dots k}=W_{ij}^{\text{RND}}+\zeta I_{ij\dots k}+\rho W_{ij}^{\text{Hop}},$ are written as
@@
Figure 4 presents a schematic phase diagram in the $\zeta$-$\rho$ plane for a maximally fluid system – one with fully antisymmetric hidden interactions ($\phi=90^{\circ}$) – in the asymptotic regime (at the $100^{th}$ time step). Whatever the strength of the implications, there is clearly a cross-over from a fluid to a stable regime once the Hebbian term becomes effective, $\rho\gtrsim 1$: the Hebbian term then stabilizes the system even when the other interactions are strictly antisymmetric.
Figure 5 shows a similar phase diagram in the $\phi$-$\rho$ plane, when implications are negligible, $\zeta=0.01$ (note that $\rho$ has been moved to the y-axis). It appears to demonstrate that the stable regime, when syntactic parameters have completely stopped evolving by the $100^{th}$ time step, can be realized with either of two conditions, or both: when the Hebbian term is strong, $\rho>1$, or hidden interactions are sufficiently symmetric, $\phi<30^{\circ}$. The regime boundaries appear surprisingly close to a Boolean rule.
@@
Figure 6: Different languages rapidly converge, for $\rho>1$. Mean pairwise Hamming distances between realizations of distinct languages subject to identical hidden couplings, after only two time steps (($t=2$, panel (a)), and in the asymptotic regime ($t=100$, panel (b)). These distances vanish when the Hebbian term is dominant ($\rho>1$), indicating rapid collapse, already after 2 time steps, onto a single syntactic structure. Note that in the fluid regime, bottom right corners, Hamming distances $\sim 47$ imply that typically half the parameters agree, between any two vectors, and half do not.
In the remainder of this study, therefore, the Hebbian term will not be considered further, aside from a detail in Fig. 9.
@@
Figure 8: Evolution of inter-language and intra-language distances. The diagrams, in the $\zeta$$\phi$ plane (at $\rho=0$), show both at early ($t=2$, top row) and asymptotic times ($t=100$, bottom row) the average Hamming distance between simulations starting from the same natural language (distance in condition S, left) and different natural languages (distance in condition D, middle), as well as their difference, (D-S, right), when using always identical hidden couplings. With time, both average distances diminish in areas characterized by strong implicational dynamics, especially when asymmetry is minimal. Still, they do not vanish: the average Hamming distance D between descendants of different languages tends (asymptotically, or for $t=100$) to decrease from D $\sim 35-37$ for weak implications ($\zeta<1$) thorough an intermediate range (e.g., D $\sim 18$ for $\zeta=10$, the star symbol) down to values D $\sim 3.1$ when implications dominate, for $\zeta>100$ (round symbol). Their average difference D-S illustrates the gradual fading of the phylogenetic/historical signal, i.e., the trace of shared ancestry among languages, everywhere except in the nearly symmetric and low (or at most intermediate) implicational strength regime, bottom left. Even there, the signal is however weak, a difference in at most 454-5 parameters, compared to the dispersion observed starting from the same initial condition (left column).
@@ V Preservation of a phylogenetic signal
As an example of temporal evolution of a syntactic system within our model, we compare the presumed asymptotic state at $t=100$ with selected intermediate stages ($t=2,10,40$) for the case of a real-world starting point — Italian — in Fig. 7. When implications are irrelevant and asymmetry is high, the vector at $t=100$ remains very distant from its ancestors at essentially all preceding $t$ stages – note that a distance $\sim 47$ implies that half the 94 parameters take the same value in the two configurations, half a different value: effectively no syntax correlation. In other words, parameter values wander randomly in syntactic space, and reach no stable state nor do they approach an asymptotic state. Something similar occurs, albeit with reduced fluidity, when implications are stronger, but hidden couplings are still asymmetric. In contrast, the distance from the asymptotic configuration gradually vanishes if hidden couplings are prevailingly symmetric, and in fact when they are close to full symmetry syntactic parameters are already stable after a few time steps. These results are consistent across different languages.
@@
If that is the time course of simulations starting from the same point, we need to compare them with those starting from other initial syntax vectors. A phylogenetic signal encodes the relatedness among descendants of the same syntactic configuration that persists, despite the initially rambling trajectories in syntactic space, produced by the hidden couplings. We obviously need to keep the assignment of hidden couplings the same, and compare distances between different runs with the Same starting point (case S), and with Different starting points (case D). Figure 8 presents this comparison, as a grand average, in the entire $\zeta$-$\phi$ plane (at $\rho=0$), for configurations evaluated at two distinct time points: $t=2$ (top row) and $t=100$ (bottom row), corresponding to the early and asymptotic phases of the model, respectively. In each case, the main factor affecting these distances, which we label simply S and D like the cases they refer to, is the strength of the implications: when they are strong, they constrain the dynamics and both types of distance grow less - in fact, they decrease, reaching lower values at $t=100$ than at $t=2$. Thus, while for minimal asymmetry and in the low implication region, $\zeta<1$, the average D $\gtrsim 35$, in the strong implication regime, $\zeta>100$, it goes down below one tenth of this value, D $\sim 3.1$. When implications compete with the hidden couplings, $1<\zeta<100$, one observes intermediate values; but, whatever $\zeta$, there is always dispersal, i.e., a multiplicity of trajectories, which do not converge to a single language.
@@
Interestingly, however, when we take the difference S-D, the only region where it remains above zero at $t=100$ (even though with small values) is the region of nearly symmetric hidden couplings and weak or intermediate implications, up to about $\zeta\approx 50$. Strong implications, $\zeta>100$, even with symmetric hidden couplings, retain a phylogenetic signal only for the first few time steps (top left region in the top right panel of Fig. 8). This indicates that in this first, simply binary model the phylogenetic signal is washed away, after a sufficient number of time steps, if implications are strong.
@@
Figure 9: Temporal dynamics of the phylogenetic signal. (a) The time course of the phylogenetic signal with $\phi$ fixed at $15^{\circ}$ and $\rho=0$, showing the rapid decrease in signal followed by stabilization at a steady-state value. A robust asymptotic signal is still observed for $\zeta=1$, i.e., when implications and hidden contributions would be balanced, except that implications are much sparser than the all-to-all hidden connectivity. As $\zeta$ increases, indicating a greater influence of implications over hidden terms, the steady-state signal decreases. In the refined model with the 0 state, however, the total signal (dashed blue curve) stabilizes at a level similar to that of the binary model with $\zeta=10$. (b) The impact of varying $\rho$ and $\phi$, while keeping $\zeta=10$ constant. With limited asymmetry, $\phi=15^{\circ}$, a small Hebbian term (the red dashed line for $\rho=0.1$) leads the system to approximate the steady-state signal faster (compare to the yellow line for $\rho=0$). Further asymmetry, i.e., increasing $\phi$ leads to the vanishing of the signal, concurrent with the cross over to the fluid regime. (c) The dependence on $\phi$ of the steady state signal (after 100 time steps) is similar, though not identical, for the model with the 0 state (total signal, dashed blue curve) as for the binary model (red curve).
@@
VI The model with undefined parameters
In the version of the model considered so far, implications are treated as contributing to the weighted input sum that determines the value of the affected parameter; they are strictly asymmetric and can involve multiple influences, but are otherwise on par with the hidden couplings. By decreasing their strength $\zeta$ they can be made into soft constraints or even slight biases, which enables the correspondence with the simple spin glass model in which they do not appear at all. It is possible, however, to implement them as precise, hard constraints that operate before ever taking the weighted sums. To that end, first we define a value 0 for the undefined state of parameters, distinct from their default value 1-1.
Figure 10: Fluidity in the model with the 0 state. The number of parameters changing value at time step 100 is plotted (above) as a function of the asymmetry angle $\phi$, separately for the parameters that enter or leave the 0 state, and for those flipping between values $\pm 1$, and compared with those flipping in the binary model with $\zeta=1000$. The bars (below) represent the same data – the sum of both types of change in the case of the revised model – with the color coding of Fig.3, and include results obtained with 3 non-zero $T$ values. The ‘hard constraints’ model presents a similar $\phi$-dependence, for $T=0$, as the binary model with $\zeta=1000$, unlike the fluidity of the binary model with $\zeta\to\infty$. With higher $T$, the fluidity is gradually reinstated also for small $\phi$, but the model with $T\ll 1$ behaves similarly to the $T=0$ one: for $\phi<30^{\circ}$ both types of change scale $\propto T$.
Partitioning parameters not in the classes $D_{k}$ mentioned above, but rather in levels (those at level $L_{k}$ are constrained by implications involving at least one parameter from level $L_{k-1}$, and in the PCM dataset $k=0,\dots,10$, see Fig.A.1), we then update their values going down the hierarchy, from $L_{0}$ to $L_{10}$, within each time step. For each level, each time step uses a different permutation of the parameters at that level, and considers them sequentially. First, it is checked whether the parameter is defined, according to the implications and to the values of the parameters above; if it is defined, it is updated using an equation with only the hidden couplings, setting it to the value +1 with probability
$P(S^{\mu}_{i_{p}}(t+1)=1)=\frac{1}{1+e^{-\left[\sum_{j}W_{ij}S^{\mu}_{j}(t^{\prime})-\theta_{i}\right]/T}}$ (4)
which when the ‘effective temperature’ $T=0$ reduces to Eq.(3) but without implications (or $\zeta=0$). When updating the value of one parameter at time step $t+1$, again we take the updated value at $t+1$ for all parameters above and those already updated at the same level, and the value at $t$ for those below and those at the same level not yet updated. Note that parameters in the undefined state (at $t^{\prime}$) are set to the value 0 and do not contribute to the hidden interactions. The strength of the implications $\zeta$ becomes irrelevant: they are now effectively of infinite strength. The thresholds $\theta_{i}$ remain, but acquire a different function: they can now be used to bias the setting of the marked value +1, so that the probability with which it is taken by a given parameter approximates the frequency with which it is observed across natural languages.
With these provisions, the dynamical behavior of the refined model remains similar to that of its simpler binary counterpart. This can be seen in Fig.10: at $T=0$ the parameter network is trapped in glassy states for $\phi<30^{\circ}$, like the binary model with finite $\zeta$. Increasing the ‘temperature’ leads to fluid dynamics for all $\phi$ values, but gradually (the mean number of flips can be seen to be $\propto T$), unlike the sudden change of behaviour occurring when, in the binary model, the hidden couplings become strictly negligible with respect to the implications – the $\zeta\to\infty$ limit. The model with the 0 state is thus robust, i.e., its low-temperature dynamics is still glassy.
VII Conclusions and future perspectives
@@
This study presents a framework for modeling syntactic evolution through the lens of disordered systems, bridging formal linguistics and statistical physics. By treating syntactic parameters as binary variables subject to both logical and typological implications and to hidden interactions, the model aims to capture the dual behavior of syntactic systems: their remarkable persistence across time and their gradual yet unpredictable diversification.
@@
Our results indeed indicate that in a portion of its phase space the network exhibits glassy dynamics, slow and gradually diverging, similar to that of spin glasses. Whether in the $N\to\infty$ limit the glassy regime would become a bona fide spin glass phase remains to be seen, but the glassy behavior is evident already in the network of limited size that models the $N=94$ syntactic parameters of the PCM database. Thus, the model suggests that syntactic evolution may not merely be the product of determinable, or even functionally driven processes, and might instead be shaped by subtle, usually unobservable interactions that may likely arise, e.g., from the way syntax is realized in the brain of the speakers. For $0<\phi<\phi_{c}=30^{\circ}$ the isolated model would get stuck into a stable state, which would be only metastable, that is, slowly changing, in the realistic scenario in which external stimuli are taken into consideration; for $\phi>\phi_{c}=30^{\circ}$, instead, syntax would continue to rapidly diversify, to the extent of losing any signal of persistence or indeed of realizing any shared syntactic rules even within the same community of speakers. The critical asymmetry value $\phi_{c}=30^{\circ}$, presumably marking a phase transition in the parallel statistical physics model, can serve as a valid reference point for the syntax model as well, as it is not much affected by other factors, in particular how the implications are modeled and, if they are taken as soft constraints, by their strength $\zeta$. Remarkably, the same value appears to hold if implications are taken to be hard constraints, in which case $\zeta$ becomes irrelevant. The suggestion from both versions of the syntax model, then, is that natural language parameters interact, perhaps due to their implementation in the brain, via couplings that are, if not strictly symmetric, at least prevailingly symmetric (quantitatively, $\phi<30^{\circ}$).
@@
Of course, in our idealized framework – no external inputs whatsoever and zero or very low ‘temperature’, i.e., minimal noise – syntactic systems can become ‘frozen’ in certain configurations, while in more realistic scenarios they would continue to evolve, though slowly, in the corresponding region of the phase diagram. It should be noted that a separate study has shown that a disordered network of binary variables interacting randomly with each other as well as with a set of multi-state variables (also randomly) can be even glassier, that is slower, than if it were isolated [ryom2023speed]. The suggestion would then be that if syntax – the binary parameters – interacts randomly with e.g. the lexicon, or phonology, or semantics – which can all be formulated in terms of multi-state variables – it might evolve even slower than on its own.
@@
The introduction of a Hebbian memory term would stabilize syntactic configurations, even with asymmetric hidden couplings. However, an effective Hebbian term ($\rho>1$) would lead all languages to collapse onto the same syntax, at least in our implementation of the Hebbian term. More sophisticated versions of the same idea might avoid the collapse. For example, it might be argued that treating closely related languages, like several of those from the Indo-European family that comprises nearly half of the current PCM database, as adding their separate Hebbian contributions might be excessive, theoretically dubious, and might be the key factor leading to the collapse. Still, it is not obvious at this stage how to construct a ‘balanced’ Hebbian term, mimicking the effect of orthogonal memories in autoassociative memory networks [hopfield1982neural].
We have seen that in the first model, even in the absence of a Hebbian term, the phylogenetic signal rapidly vanishes if implications are sufficiently strong to really act as hard constraints, $\zeta\gtrsim 10-50$ (see Fig. 8); this does not happen however in the more refined model with the zero or undefined state. Therefore, we can conclude that indeed adding weak hidden interactions $W_{ij}$ to strong explicit implications suffices to lead to both divergence (the multiplicity of metastable states) and persistence (the phylogenetic signal).
This study, in summary, aims to present the first steps towards mapping syntactic parameters onto a recurrent network model, and at a rather general level it can be said that model is compatible with empirical observations, i.e., it fits the data. Can it also lead to predict more detailed data, yet to be collected? Clearly, before being able to make genuine predictions, the model needs to be developed in several directions. We mention here a few of them.
Two developments that are conceptually trivial, but might have a major practical effect, and which involve a lot of hard work, would be to extend the set of parameters and the set of languages now included in the PCM database. Current versions of the database include only Eurasian languages, and among them an over-representation of Indo-European ones. It is often assumed, but not really known, that the main statistical properties of their distribution of parameter values reflect more general trends among all world languages. Similarly, the database is restricted to parameters affecting the determiner phrase, perhaps half or less of those relevant to syntax as a whole. Would including other parameters bring about differences in model dynamics? Along these two important directions, progress will be possible once a significantly extended PCM database is made available.
Progress that depends on computational modeling work, instead, involves replacing the random assignment of hidden weights with an educated guess based on the available statistics of parameter values. This is called ‘network reconstruction’, and much effort has been invested in developing it, for example in the context of protein folding [morcos2011direct] or recording the activity of multiple single neurons [roudi2009ising]. Appropriate methods are still to be explored in the case of syntax, which would seem very straightforward, were it not for the implications. With the random assignment, natural languages turn out to be far from any of the multiple quasi-stable configurations of parameter values the network rapidly approaches in the first few time steps, as those configurations depend on hidden weights that bear no relation to existing languages. This prevents making any form of detailed predictions about the evolution of the syntax of individual languages. The hope, then, is that effective strategies for network reconstruction will eventually lead to specific predictions or accounts of specific instances of diachronic change.
An even more ambitious goal is to make inferences about parameter couplings from electrophysiological (EEG/MEG) data, neuropsychological observations in patients with language impairments or studies of language acquisition in children. While the first two seem remote possibilities at the moment, it is encouraging to see the beginning of a structured understanding of the sequential processes of parameter acquisition [friedmann2021growing] (but see [bosch2025another]). The emerging sequential patterns, once generalized across a significant number of languages, represent a form of partial ordering that can be combined with the ordering associated with implications, as in Fig.A.1; further, parameters should be distinguished depending on the type of their putative default state [longobardi2023grammatical, crisma2025your] and on the ease with which each parameter can be set, by a speaker acquiring the language, from a corpus of spoken utterances [crisma2025your]. These different lines of evidence can aid the reconstruction of a less random assignment of hidden weights, to ultimately bring the model closer to real language phenomenology, as we plan to pursue in future work.
@@
Acknowledgments. We are indebted to the many researchers who have contributed to the PCM database, foremost Cristina Guardiano and Paola Crisma; and to Eric De Giuli for stimulating discussions.
@@
Data Availability. The PCM Database has been published, e.g. in [ceolin2021boundaries] and [crisma2025your] and expanded versions are likely to be published in the near future. The software used in the present study will be made freely available upon publication.
@@
Author Contributions. H.Y. carried out the numerical experiments and drafted the manuscript; K.I.R. contributed to the model formulation and carried out early simulations; G.L. conceptualized the study, structured the syntactic implications in the PCM database and revised the manuscript; A.T. conceptualized the study, coded the implications as polysynaptic weights and drafted the manuscript.
Appendix A Syntactic dependence